sec/ohm is equal to


1
Expert's answer
2022-05-27T09:38:54-0400
"\rm\frac{s}{\Omega}=\frac{s}{\frac{kg\cdot m^2}{s^3\cdot A^2}}=\frac{s^4\cdot A^2}{kg\cdot m^2}"

"\rm\frac{s}{\Omega}=\frac{s}{s/F}=F"